A mix of O, N, and Ar has total pressure of 1.5 atm. The weights of the gases in the mix are 1.42g, 1.00g, and .50g. Calculate partial pressure of each gas; if volume of mix is 2L, what's the temp of the gas? At this temp, what's the root mean square speed of the N molecules?
1
Expert's answer
2012-12-04T06:30:21-0500
Pi =xi*P
xi = ni/n
n(O2) = 1.42 g / 32 g/mol = 0.044 mol
n(N2) = 1.00 g / 28 g/mol = 0.0357 mol
n(Ar) = 50 g / 40 g/mol = 1.25 mol
xi(O2) = 0.044/(0.044 + 0.0357+1.25) = 0.044/1.33 = 0.033
xi(N2) = 0.0357/1.33 = 0.027
xi (Ar) = 1.25 / 1.33 =0.94
Pi(O2) = 0.033*1.5= 0.0495 atm
Pi(N2) = 0.027 * 1.5 = 0.0405 atm
P (Ar) = 0.94*1.5 = 1.41 atm
PV = nRT
T = PV/nR
T = 15450 Pa * 0.002 m3 / 1.33 mol *8.314 = 2.79 K